# Touch in the Wild: Learning Fine-Grained Manipulation with a Portable Visuo-Tactile Gripper

### v1

@@ Abstract
Figure 1: (a) Our portable handheld gripper enables synchronized collection of visual and tactile data through integrated multimodal sensing, supporting large-scale data collection in the wild. (b) We introduce a multimodal representation learning framework that fuses visual and tactile inputs to support fine-grained downstream manipulation tasks.
@@ 1 Introduction
Two key challenges have prevented widespread visuo-tactile data collection in the wild: (i) Portable tactile hardware. Most existing tactile sensors are bulky, rigid, or not robust enough. For example, soft bubble sensors Kuppuswamy et al. (2020) are too large for handheld use, while other optical-based sensors Yuan et al. (2017) are often bulky or less robust for long-term in-the-wild use, making them less suitable for mobile or outdoor deployment. (ii) Learning from multimodal data. Tactile and visual signals differ significantly in scale and in the nature of collected information. Tactile inputs are local and physical, while vision encodes broader spatial context. Learning effective representations that integrate both modalities–particularly from large-scale, unstructured datasets–remains an open challenge.
@@ 2 Related Works
Scalable multi-sensory data and learning. Reinforcement and imitation learning have become central to robotic manipulation Yin et al. (2023); Huang et al. (2023); Su et al. (2024); Wang et al. (2024); Ha et al. (2024); Chi et al. (2024); van Hoof et al. (2016), but progress is often bottlenecked by the lack of tactile sensing or access to large-scale, high-quality visuo-tactile datasets. While simulation can mitigate data scarcity, simulated tactile signals tend to diverge significantly from their real-world counterparts, limiting transferability Akinola et al. (2024); Su et al. (2024); Wang et al. (2022); Si and Yuan (2022). This makes scalable real-world visuo-tactile data collection increasingly important. However, acquiring multi-sensory data at scale remains challenging. Beyond RGB streaming, each additional modality Gao et al. (2022); Yang et al. (2024); Dou et al. (2024); Yang et al. (2023); Li et al. (2022)–such as audio Liu et al. (2024), force, or tactile–introduces added hardware complexity, synchronization overhead, and environmental constraints. As a result, most prior efforts have limited data collection to structured indoor environments Huang et al. (2024).
@@
However, tactile sensing is most valuable in uncontrolled, in-the-wild settings, where vision may degrade due to poor lighting or background clutter Wang et al. (2023); Finn et al. (2017), while contact forces remain stable. Prior “in-the-wild” systems have demonstrated strong vision-only performance on simple tasks, but they overlook the complementary benefits of touch Chi et al. (2024); Wu et al. (2024). To bridge this gap, we propose a portable handheld visuo-tactile system that combines a fisheye RGB camera with a lightweight, flexible tactile array. This setup enables synchronized, large-scale collection of visual and tactile data across diverse, unstructured environments. Leveraging this system, we build a visuo-tactile dataset and train a masked reconstruction encoder that (i) accurately recovers tactile signals, and (ii) enhances downstream policy learning through visuo-tactile representations. In contrast to previous methods that directly merge vision and touch in a unified 3D space Huang et al. (2024), our approach enables scalable cross-modal representation learning, fostering stronger correlations between visual and tactile modalities for fine-grained manipulation.
@@
Visuo-tactile manipulation. Tactile feedback plays a critical role in human manipulation, particularly when visual input is occluded or ambiguous Johansson and Westling (2004); Jenmalm and Johansson (1997). Similarly, tactile sensing enables more adaptive robotic manipulation for tasks that are complex or require precision. As a result, there is growing interest in integrating vision and touch to improve robotic manipulation Huang et al. (2024); Suresh et al. (2023); Ding et al. (2023); Lee et al. (2019); van Hoof et al. (2016); Chen et al. (2022); Hansen et al. (2022); Yuan et al. (2023); Qin et al. (2023); Chen et al. (2022); George et al. (2024); Yang et al. (2022); Dave et al. (2024); Huang et al. (2025). Much of this prior work relies on optical tactile sensors, which image surface deformations to infer contact geometry and texture Calandra et al. (2017); Yuan et al. (2018, 2023); Kolamuri et al. (2021); Calandra et al. (2018); Xue et al. (2025); Dou et al. (2024). While these sensors provide rich signals, they are typically rigid, bulky, and less robust for long-term in-the-wild use, limiting their applicability in portable or unstructured settings.
@@
In contrast, we use thin, flexible tactile sensors Huang et al. (2024); Sundaram et al. (2019); Feng et al. (2025) that directly measure force distributions over the contact surface. Embedded in soft robotic fingers, these sensors provide consistent, object-agnostic representations that are easier to generalize and better suited for scalable, real-world learning. Although such sensors have been explored in lab setting Huang et al. (2024); Bicchi and Kumar (2000), large-scale, in-the-wild visuo-tactile datasets with synchronized RGB and tactile data capturing physical interactions have not previously been available. Our work fills this gap by collecting–and publicly releasing–a diverse, real-world visuo-tactile dataset. It spans a wide range of tasks and environments, laying the foundation for scalable multimodal learning and robust policy development. We acknowledge the concurrent work ViTaMIn Liu et al. (2025), which also proposes a portable visuo-tactile data collection system and a cross-modal learning framework for contact-rich manipulation. Both of our work shares high-level goals, but the system design, representation learning approach, and task emphasis differ in several key aspects. We appreciate their contribution to this emerging area and encourage the reader to refer to their work for complementary insights.
@@ 3.1 Scalable Flexible Tactile Sensors
As shown in Fig. 2 (a), we embed thin, matrix-based tactile pads into the soft, fin-shaped fingers of our handheld gripper. The sensor architecture builds on the triple-layer design from 3D-ViTac Huang et al. (2024), adapted to fit the geometry of the adaptive fin-shaped gripper Chi et al. (2024). Each tactile pad consists of a piezoresistive sensing layer sandwiched between two flexible printed circuits (FPCs). To accommodate the elongated, flexible fins, we introduce two key modifications: (1) Higher spatial resolution. Capturing contact patterns along the finger’s length requires denser spatial sampling. The stainless-steel electrodes used in 3D-ViTac limit both resolution and signal stability. By replacing them with FPC electrodes, we achieve uniform trace pitch, improved robustness, and a per-pad resolution of $12\times 32$ taxels, each measuring a $2\times 2mm^{2}$ area. This allows us to capture fine-grained, dynamic contact signals. (2) Rapid, scalable fabrication. The use of FPCs enables tool-free assembly. Each pad can be fabricated in under five minutes and mounted on the gripper in an additional two, supporting scalable deployment for large-scale tactile data collection.
@@ 3.2 Portable Multi-Modal Sensing System
At the firmware level, we optimize the serial protocol to stream each $12\times 32$ pad at $23$ Hz and synchronize with the visual information from the fish-eye camera. Tactile frames are timestamped directly on the microcontroller and transmitted over USB to a host device (e.g., a laptop or any portable Ubuntu system). The battery-powered, handbag-sized system is easily deployed in grocery stores, outdoor markets, and other unstructured, in-the-wild environments (Fig. 2 (a)), enabling high-throughput and scalable visuo-tactile data collection.
@@ 4 Visuo-Tactile Representation and Policy Learning
To perform precise manipulation, robots must effectively integrate both visual and tactile signals. RGB images provide global, semantic context—such as object identity and workspace layout—while tactile signals offer local, contact-rich feedback that is often occluded from vision Li et al. (2019). These complementary modalities follow different statistical distributions, making it non-trivial to learn unified representations that preserve modality-specific information while enabling effective cross-modal reasoning. We propose a two-stage learning framework (Fig. 3) that first learns a joint visuo-tactile representation via masked tactile reconstruction, and then integrates the learned representation into a diffusion policy Chi et al. (2024) for downstream manipulation tasks.
@@ 4.2 Stage 1: Visuo-Tactile Representation Learning
While prior work often relies on contrastive learning to align embeddings from different modalities Dave et al. (2024); Guzey et al. (2023); Liu et al. (2025), such objectives tend to suppress the fine-grained, geometry-sensitive signals captured by tactile sensors. Instead, we adopt a masked autoencoding objective He et al. (2022), which reconstructs missing tactile regions conditioned on partially observed tactile input and visual context. This formulation encourages the encoder to retain tactile-specific information while leveraging vision for inference.
@@
Tactile encoder. Each tactile reading consists of two fingertip arrays, each of shape $1\times 12\times 32$, stacked vertically to form a $1\times 24\times 32$ tactile image. We then apply a fixed colormap to convert this single-channel map into a 3-channel RGB tactile image. This image is divided into non-overlapping $4\times 4$ patches, resulting in a $6\times 8$ patch grid. During training, we randomly mask 60–80% of the patches in 95% of samples using a learnable token $T_{\mathrm{mask}}$; the remaining 5% are shown in full. The masked tactile input is defined as:
$T_{\mathrm{visible}}=M\odot T+(1-M)\odot T_{\mathrm{mask}},$ (2)
where $M\in\{0,1\}^{6\times 8}$ is a binary patch mask. The visible tactile input $T_{\mathrm{visible}}$ is processed by a 3-layer CNN to produce a $d$-dimensional embedding ${z}_{\mathrm{tac}}$, where $d=768$.
Vision encoder. The RGB image $I$ is processed by a ViT-B/16 encoder initialized from CLIP Radford et al. (2021). We finetune all layers with a learning rate of $3\times 10^{-5}$, and extract the final [CLS] token as the $d$-dimensional visual embedding ${z}_{\mathrm{img}}$.
@@
[…]
@@
Tactile reconstruction decoder. The fused feature ${z}_{\mathrm{fusion}}$ is passed through a two-layer MLP followed by a sigmoid activation to produce the reconstructed tactile image $\hat{T}\in\mathbb{R}^{1\times 24\times 32}$, i.e.,
$\hat{T}=\mathrm{D}_{\psi}({z}_{\mathrm{fusion}})$
We use a full-image reconstruction loss:
$\mathcal{L}_{\text{stage1}}(\phi,\psi)=\bigl{\|}T-\hat{T}\bigr{\|}_{2}^{2},$
which encourages both local contact inference and global structural understanding.
@@ 4.3 Stage 2: Policy Learning via Behavior Cloning
Observation space. At each timestep $t$, the robot first receives raw sensory inputs
$(I_{t},\,T_{t},\,p_{t}),$
where $I_{t}$ and $T_{t}$ are the RGB and tactile images, and $p_{t}$ denotes the proprioceptive state (e.g., end-effector pose, gripper width). $I_{t}$ and $T_{t}$ are then passed through the pretrained encoder to produce the visuo-tactile embedding $z_{t}=E_{\phi}(I_{t},T_{t})$. The diffusion policy is then conditioned on
$o_{t}=(z_{t},\,p_{t})\,,$
i.e. the fused visuo-tactile embedding together with proprioception.
Diffusion policy. We adopt a conditional diffusion policy Chi et al. (2024). Instead of regressing actions directly, we train a noise predictor:
$\hat{\epsilon}^{k}_{t}=\epsilon_{\theta}\bigl{(}a^{k}_{t},\;o_{t},\;k\bigr{)},$ (6)
where $a^{k}_{t}$ is the noisy action at diffusion step $k$, $o_{t}$ is the full conditioning observations, and $\hat{\epsilon}^{k}_{t}$ is the predicted noise. During training, we sample $k$ uniformly and add Gaussian noise $\epsilon^{k}_{t}$ to the ground-truth action $a^{0}_{t}$. We formulate our loss as the MSE loss between added and predicted noise:
$\mathcal{L}_{\text{stage2}}=\mathbb{E}_{t,k}\Bigl{[}\bigl{\|}\epsilon^{k}_{t}-\hat{\epsilon}^{k}_{t}\bigr{\|}_{2}^{2}\Bigr{]}.$ (7)
At inference, we start from $a^{K}_{t}\!\sim\!\mathcal{N}(0,I)$ and denoise for $K$ steps:
$a^{k-1}_{t}=\alpha\bigl{(}a^{k}_{t}-\gamma\,\epsilon_{\theta}(a^{k}_{t},o_{t},k)\bigr{)}\;+\;\mathcal{N}(0,\sigma^{2}I),$ (8)
@@
Training details. We use Diffusion Policy’s convolutional U-Net Ronneberger et al. (2015) with DDIM inference Song et al. (2020). The model is conditioned on the fused visuo-tactile embedding and two consecutive proprioceptive observations. All encoder components—including the CLIP backbone, tactile CNN, and cross-attention layers—are finetuned during this stage using a learning rate of $3\times 10^{-5}$.
@@ 5 Experiments
In this section, we address several key questions regarding the role of touch in fine-grained manipulation and the impact of different pretraining strategies on downstream tasks. Specifically, we investigate: (1) How does touch enable fine-grained manipulation? (2) How does pretrained visuo-tactile encoder aid robust policy learning? (3) How do variations in pretraining affect downstream task performance?

### v2

@@ Abstract
Figure 1: (a) Our portable handheld gripper enables synchronized collection of visual and tactile data, supporting large-scale data collection in the wild. (b) We introduce a multimodal representation learning framework that fuses visual and tactile inputs to support fine-grained downstream manipulation tasks.
@@ 1 Introduction
Two key challenges have prevented widespread visuo-tactile data collection in the wild: (i) Portable tactile hardware. Most existing tactile sensors are bulky, rigid, or not robust enough. For example, soft bubble sensors kuppuswamy2020soft are too large for handheld use, while other optical-based sensors yuan2017gelsight are often bulky or less robust for long-term in-the-wild use, making them less suitable for mobile or outdoor deployment. (ii) Learning from multimodal data. Tactile and visual signals differ significantly in scale and in the nature of collected information. Tactile inputs are local and physical, while vision encodes broader spatial context. Learning effective representations that integrate both modalities–particularly from large-scale, unstructured datasets–remains an open challenge.
@@ 2 Related Works
Scalable multi-sensory data and learning. Reinforcement and imitation learning have driven major advances in robotic manipulation yin2023rotating; huang2023dynamic; su2024sim2real; wang2023gendp; ha2024umilegs; chi2024diffusionpolicy; hoof2016stable, yet progress remains limited by the lack of tactile sensing and large-scale visuo-tactile datasets. Simulation can partially address data scarcity, but simulated tactile signals often diverge from real-world contact dynamics, reducing transfer effectiveness akinola2024tacsl; su2024sim2real; wang2022tacto; si2022taxim. This makes scalable real-world visuo-tactile data collection increasingly important. Acquiring multi-sensory data at scale, however, is difficult. Each additional modality—beyond RGB video—adds hardware complexity, synchronization overhead, and environmental constraints gao2022objectfolder; yang2024binding; dou2024tactile; yang2023generating; li2022see; liu2024maniwav. Consequently, most existing datasets are confined to structured indoor environments huang20243d. To address this, recent work has shifted focus toward large-scale visuo-tactile pretraining, which leverages existing data to learn shared representations that generalize across sensors and manipulation tasks higuera2024sparsh; zhao2024transferable. These approaches demonstrate that pretraining can reduce task-specific data demands and better align visual and tactile modalities, laying the foundation for scalable multimodal learning.
@@
However, tactile sensing is most valuable in uncontrolled, in-the-wild settings, where vision may degrade due to poor lighting or background clutter wang2023mimicplay; finn2017one, while contact forces remain stable. Prior “in-the-wild” systems have demonstrated strong vision-only performance on simple tasks, but they overlook the complementary benefits of touch chi2024universal; wu2024fast. To bridge this gap, we propose a portable handheld visuo-tactile system that combines a fisheye RGB camera with a lightweight, flexible tactile array. This setup enables synchronized, large-scale collection of visual and tactile data across diverse, unstructured environments. Leveraging this system, we build a visuo-tactile dataset and train a masked reconstruction encoder that (i) accurately recovers tactile signals, and (ii) enhances downstream policy learning through visuo-tactile representations. In contrast to previous methods that directly merge vision and touch in a unified 3D space huang20243d, our approach enables scalable cross-modal representation learning, fostering stronger correlations between visual and tactile modalities for fine-grained manipulation.
@@
Visuo-tactile manipulation. Tactile feedback is essential for human manipulation, especially when vision is occluded or ambiguous Johansson2004RolesOG; Jenmalm4486. Likewise, tactile sensing enables robots to perform more adaptive and precise manipulation. Consequently, there is growing interest in combining vision and touch to enhance robotic manipulation huang20243d; suresh2023neural; 10.3389/fnbot.2023.1181383; lee2019making; hoof2016stable; chen2022visuo; hansen2022visuo; yuan2023robot; qin2023dexpoint; george2024vital; yang2022touch; mvitac; huang2025vtrefine. Much of this prior work relies on optical tactile sensors, which image surface deformations to infer contact geometry and texture calandra2017feeling; yuan2018active; yuan2023robot; kolamuri2021improving; calandra2018more; xue2025reactive; dou2024tactile. While these sensors provide rich signals, they are typically rigid, bulky, and less robust for long-term in-the-wild use, limiting their applicability in portable or unstructured settings.
@@
In contrast, we use thin, flexible tactile sensors huang20243d; Sundaram2019LearningTS; feng2025anytouch that directly measure force distributions over the contact surface. Embedded in soft robotic fingers, these sensors provide consistent, object-agnostic representations that are easier to generalize and better suited for scalable, real-world learning. Although such sensors have been explored in lab setting huang20243d; 844081, large-scale, in-the-wild visuo-tactile datasets with synchronized RGB and tactile data capturing physical interactions have not previously been available. Our work fills this gap by collecting–and publicly releasing–a diverse, real-world visuo-tactile dataset. It spans a wide range of tasks and environments, laying the foundation for scalable multimodal learning and robust policy development. We acknowledge the concurrent project ViTaMIn liu2025vitamin, which introduces a portable visuo-tactile data collection system and a cross-modal learning framework for contact-rich manipulation. While our efforts share similar goals, our system design, representation learning strategy, and task focus differ in several key aspects. We appreciate their contribution and encourage readers to consult their work for complementary insights.
@@ 3.1 Scalable Flexible Tactile Sensors
As shown in Fig. 2 (a), we embed thin, matrix-based tactile pads into the soft, fin-shaped fingers of our handheld gripper. The sensor architecture builds on the triple-layer design from 3D-ViTac huang20243d, adapted to fit the geometry of the adaptive fin‑shaped gripper chi2024universal. Each tactile pad consists of a piezoresistive sensing layer sandwiched between two flexible printed circuits (FPCs). To accommodate the elongated, flexible fins, we introduce two key modifications: (1) Higher spatial resolution. Capturing contact patterns along the finger’s length requires denser spatial sampling. The stainless-steel electrodes used in 3D-ViTac limit both resolution and signal stability. By replacing them with FPC electrodes, we achieve uniform trace pitch, improved robustness, and a per-pad resolution of $12\times 32$ taxels, each measuring a $2\times 2mm^{2}$ area. This allows us to capture fine-grained, dynamic contact signals. (2) Rapid, scalable fabrication. The use of FPCs enables tool-free assembly. Each pad can be fabricated in under five minutes and mounted on the gripper in an additional two, supporting scalable deployment for large-scale tactile data collection.
@@ 3.2 Portable Multi-Modal Sensing System
At the firmware level, we optimize the serial protocol to stream each $12\times 32$ pad at $23$ Hz and synchronize with the visual information from the fish-eye camera. Tactile frames are timestamped directly on the microcontroller and transmitted over USB to a host device (e.g., a laptop or any portable Ubuntu system). The battery‑powered, handbag‑sized system is easily deployed in grocery stores, outdoor markets, and other unstructured, in-the-wild environments (Fig. 2 (a)), enabling high‑throughput and scalable visuo‑tactile data collection.
@@ 4 Visuo-Tactile Representation and Policy Learning
To perform precise manipulation, robots must effectively integrate both visual and tactile signals. RGB images provide global, semantic context—such as object identity and workspace layout—while tactile signals offer local, contact-rich feedback that is often occluded from vision li2019connecting. Because these modalities follow different statistical distributions, learning unified yet modality-specific representations for effective cross-modal reasoning remains challenging. We propose a two-stage learning framework (Fig. 3) that first learns a joint visuo-tactile representation via masked tactile reconstruction, and then integrates the learned representation into a diffusion policy chi2024diffusionpolicy for downstream manipulation tasks.
@@ 4.2 Stage 1: Visuo-Tactile Representation Learning
While prior work often relies on contrastive learning to align embeddings from different modalities mvitac; guzey2023dexterity; liu2025vitamin, such objectives tend to suppress the fine-grained, geometry-sensitive signals captured by tactile sensors. Instead, we adopt a masked autoencoding objective he2022masked, which reconstructs missing tactile regions conditioned on partially observed tactile input and visual context. This formulation encourages the encoder to retain tactile-specific information while leveraging vision for inference.
@@
Tactile encoder. Each tactile reading consists of two fingertip arrays, each of shape $1\times 12\times 32$, stacked vertically to form a $1\times 24\times 32$ tactile image. We then apply a fixed colormap to convert this single‐channel map into a 3-channel RGB tactile image. This image is divided into non-overlapping $4\times 4$ patches, resulting in a $6\times 8$ patch grid. During training, we randomly mask 60–80% of the patches in 95% of samples using a learnable token $T_{\mathrm{mask}}$; the remaining 5% are shown in full. The masked tactile input is processed by a 3-layer CNN to produce a $d$-dimensional embedding ${z}_{\mathrm{tac}}$, where $d=768$.
Vision encoder. The RGB image $I$ is processed by a ViT-B/16 encoder initialized from CLIP radford2021learning. We finetune all layers with a learning rate of $3\times 10^{-5}$, and extract the final [CLS] token as the $d$-dimensional visual embedding ${z}_{\mathrm{img}}$.
@@
${z}^{\prime}_{\mathrm{tac}}$ $=\mathrm{MHAttn}(Q={z}_{\mathrm{tac}},\,K={z}_{\mathrm{img}},\,V={z}_{\mathrm{img}})\xrightarrow{\text{LayerNorm}}{z}^{\prime\prime}_{\mathrm{tac}},$ (2)
${z}^{\prime}_{\mathrm{img}}$ $=\mathrm{MHAttn}(Q={z}_{\mathrm{img}},\,K={z}^{\prime\prime}_{\mathrm{tac}},\,V={z}^{\prime\prime}_{\mathrm{tac}})\xrightarrow{\text{LayerNorm}}{z}^{\prime\prime}_{\mathrm{img}}.$ (3)
@@
Tactile reconstruction decoder. The fused feature ${z}_{\mathrm{fusion}}$ is passed through a two-layer MLP followed by a sigmoid activation to produce the reconstructed tactile image $\hat{T}\in\mathbb{R}^{1\times 24\times 32}$, i.e., $\hat{T}=\mathrm{D}_{\psi}({z}_{\mathrm{fusion}})$. We use a full-image reconstruction loss $\mathcal{L}_{\text{stage1}}(\phi,\psi)=\bigl\|T-\hat{T}\bigr\|_{2}^{2}$, which encourages both local contact inference and global structural understanding.
@@ 4.3 Stage 2: Policy Learning via Behavior Cloning
Observation space. At each timestep $t$, the robot first receives raw sensory inputs $(I_{t},\,T_{t},\,p_{t})$, where $I_{t}$ and $T_{t}$ are the RGB and tactile images, and $p_{t}$ denotes the proprioceptive state (e.g., end-effector pose, gripper width). $I_{t}$ and $T_{t}$ are then passed through the pretrained encoder to produce the visuo-tactile embedding $z_{t}=E_{\phi}(I_{t},T_{t})$. The diffusion policy is then conditioned on $o_{t}=(z_{t},\,p_{t})$, i.e. the fused visuo-tactile embedding together with proprioception.
Diffusion policy. We adopt a conditional diffusion policy chi2024diffusionpolicy that predicts noise instead of regressing actions directly. The model outputs $\hat{\epsilon}^{k}_{t}=\epsilon_{\theta}(a^{k}_{t},o_{t},k)$, where $a^{k}_{t}$ is the noisy action at step $k$, $o_{t}$ the conditioning observation, and $\hat{\epsilon}^{k}_{t}$ the predicted noise. During training, $k$ is uniformly sampled and Gaussian noise $\epsilon^{k}_{t}$ is added to the ground-truth action $a^{0}_{t}$. The loss is the MSE between added and predicted noise:
$\mathcal{L}_{\text{stage2}}=\mathbb{E}_{t,k}\!\left[\|\epsilon^{k}_{t}-\hat{\epsilon}^{k}_{t}\|_{2}^{2}\right].$ (5)
At inference, actions are initialized as $a^{K}_{t}\!\sim\!\mathcal{N}(0,I)$ and iteratively denoised:
$a^{k-1}_{t}=\alpha\!\left(a^{k}_{t}-\gamma\,\epsilon_{\theta}(a^{k}_{t},o_{t},k)\right)+\mathcal{N}(0,\sigma^{2}I),$ (6)
@@
Training details. We use Diffusion Policy’s convolutional U-Net ronneberger2015u with DDIM inference song2020denoising. The model is conditioned on the fused visuo-tactile embedding and two consecutive proprioceptive observations. All encoder components—including the CLIP backbone, tactile CNN, and cross-attention layers—are finetuned during this stage using a learning rate of $3\times 10^{-5}$.
@@ 5 Experiments
In this section, we address key questions about the role of touch in fine-grained manipulation and the influence of different pretraining strategies on downstream performance. Specifically, we ask: (1) How does touch enable fine-grained manipulation? (2) How does visuo-tactile pretraining improve policy robustness? (3) How do pretraining variations affect task performance?